OpenStudy (anonymous):

On which of the following lines does the point, (1,-13) lie? a. y=5x-9 b. y=-2x+5 c. y=-5x+5 d. y=-3x-10

jimthompson5910 (jim_thompson5910):

(1,-13) means x = 1 and y = -13

jimthompson5910 (jim_thompson5910):

checking choice A y=5x-9 -13=5(1)-9 ... replace x with 1, y with -13 -13 = 5 - 9 -13 = -4 The equation -13 = -4 is false, so (1,-13) is NOT a solution to y=5x-9 the point does NOT lie on y=5x-9

jimthompson5910 (jim_thompson5910):

you will do the same for the other choices

OpenStudy (anonymous):

so i just input the values that it gives me into the equation for x and y? if that makes sense

jimthompson5910 (jim_thompson5910):

yes and you are looking for cases when you plug in (x,y) = (1,-13) and you get a true result

OpenStudy (anonymous):

i got d and it was right so thank you for helping me understand how to do that

jimthompson5910 (jim_thompson5910):

D is correct